# Construction of Functional Materials Knowledge Graph in Multidisciplinary Materials Science via Large Language Model

### v1

@@ Methods
Table 4: Human evaluation metric on randomly selected triples.
Discussion
In this research, we construct a functional material KG by fine-tuned LLMs and ensuring the traceability of all information. During the process, we encountered two main challenges and proposed corresponding solution: 1) Fine tuning LLM reduces the requirement for training set volume, but requires higher quality for each data in training set. when the quality of training dataset is limited, the performance of NER and RE is not good. Recursively generating training set data automatically from inference result is a good solution. High recall and precision data from each inference can be put into the training set after simple manual check. 2) Fine-tuned LLMs perform well on NER and RE task, but for ER task, they can not shows a same. However, the ER is also important in graph construction. To address the issue, we apply several ER methods and technologies including ChemDataExtractor, mat2Vec, Word Embedding Clustering and construct an expert dictionary storing information about energy material. Theses methods and technologies can not only achieve high perfromance ER task, but also remove some erroneous entities and relations.
To better demonstrate the reliability of our method and FMKG, we compare them with existing structural information extraction method and material KG. We employ a similar approach to NERRE task of John Dagdelen et al. [21] to extract structured information from the text. The difference is that we have applied ER method to correct entities and relations, and finish ER tasks before graph construction. We evaluated our pipeline on NERRE’s evaluation dataset. It worth to mention that cross dataset evaluation allow for flexibility of multiple plausible annotations of an entity under different labels (e.g., "cathode" in NERRE is placed in "Application" label while our expert tend to place it in "Descriptor" since "cathode" and "anode" are component of "battery" which is an application.), and presented the results in a Table 5. From the results, it is evident that our pipeline has slightly lower recall across labels compared to the NERRE task, but our method holds a competitive edge in precision. This is because we believe that the primary goal of KG construction is to ensure the accuracy of the data. Hence, in the process of ER, our strategy tends to sacrifice a portion of recall in exchange for higher precision. However, when considering both recall and precision, our method scores higher in F1 across the "Acronym", "Application", "Structure/Phase", and "Descriptor" labels compared to NERRE, with the improvement in "Acronym" being particularly significant due to its lower precision in NERRE. This underscores the advantage of our ER pipeline and indirectly attests to the credibility of the FMKG.
Label Extraction recall (NERRE) Extraction precision (NERRE) Extraction F1 (NERRE) Extraction recall (DarwinN) Extraction precision (DarwinN) Extraction F1 (DarwinN)
Formula 0.943 0.943 0.943* 0.828 1.000 0.906
Name 0.692 1.000 0.818 0.667 1.000 0.800
Acronym 0.667 0.400 0.500 0.571 1.000 0.727
Descriptor 0.576 0.905 0.704 0.781 0.962 0.862
Application 0.797 0.870 0.832 0.750 1.000 0.857
Structure/Phase 0.754 0.920 0.829 0.613 0.905 0.730
*Best F1 scores for each label.
Table 5: Evaluation comparison of NERRE and DarwinN.
Knowledge graphs specifically dedicated to general materials are rare. In our discussion, we draw a comparison with MatKG2 [15], showcasing the differences in construction methodologies between MatKG2 and FMKG, as depicted in the Figure 5. The construction of MatKG2 involves in a multi-step NER and RE pipeline, which can not catch the source of relations. In our method, we use end-to-end method that use a single LLM have been investigated for joint named entity recognition and relation extraction. It can not only catch keep source of each triple in KG so that enhance the factual basis of KG, but also perform better on RE task.
Figure 5: Schematic comparison of FMKG and MatKG2.
KGs are always dynamic, especially in cutting-edge field such as material science. The LLM internalizes a large amount of knowledge during the training process, it can perform good inference even when new nodes and relationships appear, making KG easy to update. Besides, our work is a starting point where other material researchers can continue build FMKG using our result as training dataset, using our dictionary, ChemDataExtractor and mat2vec to achieve ER task, our method can also be extend to full paper, extract more functional material information to construct a larger and more comprehensive functional material knowledge graph. Besides, our pipeline can also be used in other specialized domains to construct domain-specific KGs.
Conclusion
In this study, we introduce a new NLP pipeline for KG construction, which aim to efficiently extract the triples from unstructured scientific texts. The main feature of the method is that it can fine-tune the LLMs by annotating a small amount of data, and use the fine-tuned LLM to extract structured information from a large amount of unstructured text. The entire process does not rely on any prediction, which can maximize the authenticity and traceability of structured information. By employing this method, we construct a Functional Material Knowledge Graph (FMKG) contains the materials and their related knowledge from abstract of 150,000 peer-reviewed paper. After analyzing, we have demonstrated the effectiveness and credibility of FMKG.
In addition, our method and KG have great potential in different dimensions. Firstly, enhancing the depth of structured information extraction to encompass entire research papers promises a richer, more detailed knowledge graph. This involves not only expanding the scope of data analyzed but refining the process to capture nuances within complex scientific texts. Secondly, refining entity labels within our system allows for a more precise categorization of data, including incorporating detailed attributes such as synthesis conditions or property parameters, which significantly improves the granularity and utility of the knowledge graph. Thirdly, the versatility of our NLP pipeline suggests its applicability across different scientific domains, offering a template for constructing domain-specific knowledge graphs beyond material science. Lastly, integrating FMKG with existing knowledge graphs like MatKG opens avenues for creating a more interconnected and comprehensive dataset, facilitating advanced research and application development in material science and beyond.

### v2

@@ A. Validation for MKG
B. Predication result
As Figure 7 illustrates, we display the materials ranked in the top five for 2023 across various domains, alongside their rankings in previous years based on available data up to 2023. It is important to note that these materials-application pairs were not ranked within the top five before 2017, which is why they do not appear in the figure for those earlier years. Stars mark the years these materials were featured in literature for their respective applications. The figure emphasizes the significance and emerging utility of these materials by highlighting the top five in applications such as catalysts, batteries, fuel/solar cells, and other fields from 2017 to 2023, indicating their growing importance and potential to revolutionize technology in these areas. This predication serves as an inspiration to materials researchers, accelerating progress in material development. For example, consider the material NiFeN HC/NF; it not only excels in the Oxygen Evolution Reaction (OER) but has also demonstrated its capability as a catalyst for the Hydrogen Evolution Reaction (HER). This means that NiFeN HC/NF is a promising bifunctional electrocatalyst for water electrolysis, aligning with our understanding of its potential in dual-function applications.
C. Paramaters for LLM fine tuning
epochs: 10 train batch size (per device): 1 eval batch size (per device): 1 gradient accumulation steps: 2 evaluation strategy: "no" save strategy: "steps" save steps: 500 save total limit: 1 learning rate: 2e-5 weight decay: 0. warmup ratio: 0.03 lr scheduler type: "cosine" logging steps: 1 tf32: False fsdp: "full shard auto wrap" fsdp transformer layer cls to wrap: "LlamaDecoderLayer" token max length: 1024
Figure 7: High rank predicated materials in (a) Catalyst, (b) Fuel/Solar cell, (c) Battery and (d) Other fields, the stars represent the material has been reported in this application.
D. Experiments compute resources
The compute resources were primarily consumed in LLM fine-tuning, involving the use of 8 GPUs for 9.02 hours with a 36% utilization rate, consuming 25.98 service units and using 14245 MiB of storage from a quota of 81920 MiB.